# Performant near-term quantum combinatorial optimization

### v1

@@ Abstract
Conclusion – We developed an algorithm for quantum combinatorial optimization that makes several important advances over leading alternatives, in terms of optimization procedure and iterations, convergence behaviors, and circuit depth; widespread success of the algorithm is evident in simulations of diverse problems with up to 20 qubits. Key features of our approach include counterdiabatic Hamiltonian generators capable of preparing optimal solutions, training through QITE with a nonlinear objective that emphasizes the ground state, and resource-efficient generation of volume-law entanglement during training. We expect the low depth and minimal training requirements of our algorithm to enable quantum computers to solve combinatorial optimization problems with several tens of qubits in the immediate future, with extensions to larger sizes and scaling analyses beyond classical simulability as quantum computers continue to develop. The low depth, minimal resource requirements, high entanglement, and superior empirical performance make our approach a leading contender for near-term quantum computational advantage in combinatorial optimization.
@@
Acknowledgements – We thank Joseph Wang, Eugene Dumitrescu, Jim Ostrowski, Ryan Bennink, George Siopsis, Rizwanul Alum, Noah Bauer, and Rebekah Herrman for discussions of variational quantum optimization algorithms. We thank Maxime Dupont for discussion and sharing numerical results from Ref. Dupont and Sundar (2024). T.D.M was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center and devised the ansatz, optimization algorithm, performed the simulations, and contributed to writing. P.C.L. was supported by the DARPA ONISQ program under award W911NF-20-20051 and contributed to design, analysis, writing, and entanglement computations. This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.

### v2

@@ I Conclusion and Outlook
We have developed an algorithm for quantum combinatorial optimization that makes several important advances over leading alternatives, in terms of optimization procedure and iterations, convergence behaviors, and circuit depth. Key features of our approach include counterdiabatic Hamiltonian generators capable of preparing optimal solutions, training through VAR-IT with a nonlinear objective that emphasizes the ground state, and resource-efficient generation of volume-law entanglement during training. Widespread success of the algorithm was evident in numerical simulations of diverse problem instances with up to 20 qubits, which consistently converged to optimal solutions, as well as algorithm execution on quantum computers with up to 32 qubits.
We have demonstrated full hybrid on-QPU optimization of the MaxCut problem for graphs of varying types, connectivity and sizes. Our results have shown that a truncated version of our algorithm is capable of finding optimal solutions with 32 or fewer qubits, with convergence to high optimal solution probability. The problem instances were chosen completely at random which suggests the robustness of the algorithm in solving generic instances of the MaxCut problem, as anticipated from the numerical simulations. The high quality results also demonstrated the robustness against shot noise and QPU noise even at 32 qubits. These properties demonstrate clear advantages of our approach relative to the previous state-of-the-art.
In future work, we expect the combination of our algorithm with problem decomposition techniques Karypis and Kumar (1998a, b); Ushijima-Mwesigwa et al. (2021); Acharya et al. (2024); Ponce et al. (2023); Maciejewski et al. (2024) will enable hierarchical solutions to combinatorial problems that begin to scale towards industrially-relevant sizes. Another important frontier is the extension of our approach to higher-order binary optimization problems where classical methods fail to return optimal solutions for problems with $\lesssim 100$ variables Shaydulin et al. (2023b), where optimal solution convergence as we observed here could potentially discover better solutions than any conventional method. The low depth, minimal resource requirements, high entanglement, superior empirical performance, and scalability of our approach make it a promising candidate for solving these and other interesting problems in the future.
II Acknowledgements
@@
We thank Joseph Wang, Eugene Dumitrescu, Jim Ostrowski, Ryan Bennink, George Siopsis, Rizwanul Alam, Noah Bauer, and Rebekah Herrman for discussions of variational quantum optimization algorithms. We thank Maxime Dupont for discussion and sharing numerical results from Ref. Dupont and Sundar (2024). We thank Ruslan Shaydulin for discussion and for bringing Refs. Hao et al. (2024); Moses et al. (2023); DeCross et al. (2023); Sachdeva et al. (2024) to our attention. T.D.M was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center and devised the ansatz, optimization algorithm, performed the simulations, and contributed to writing. P.C.L. was supported by the DARPA ONISQ program under award W911NF-20-20051 and contributed to design, analysis, writing, and entanglement computations. This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Supplemental Materials: Performant near-term quantum combinatorial optimization
III Ansatz
In pursuit of a solution of the binary optimization problem, we perform an analysis of the QAOA ansatz,
$\left|\vec{\theta},\vec{\gamma}\right\rangle_{p}=\prod_{j=0}^{p}\exp(-i\gamma_{j}H_{m})\exp(-i\theta_{j}H_{c})\left|+\right\rangle^{\otimes},$ (S1)
where $H_{m}$ is the standard mixing Hamiltonian and $H_{c}$ is the cost Hamiltonian with an eigenspectrum $H_{c}\left|\bm{z}\right\rangle=C(\bm{z})\left|\bm{z}\right\rangle$ that contains the set of classical solution values
$H_{m}=-\sum_{j\in V}X_{j}$ (S2)
$H_{c}=\sum_{j,k\in E}\omega_{jk}Z_{j}Z_{k}=\sum_{\alpha}\lambda_{\alpha}P_{\alpha}\,.$ (S3)
In the limit as $p\rightarrow\infty$, and with appropriately chosen $\vec{\theta}$ and $\vec{\gamma}$, the QAOA ansatz yields adiabatic evolution from the ground state $\left|+\right\rangle^{\otimes n}$ of $H_{m}$ to the optimal-solution ground state $\left|\bm{z}_{\mathrm{opt}}\right\rangle$ of $H_{c}$. At finite $p$ that are implementable on current quantum computers, QAOA has been shown to be successful at approximating $\left|\bm{z}_{\mathrm{opt}}\right\rangle$ by variationally optimizing the parameters $p$. Analyzing the QAOA/adiabatic ansatz by combining the two sequential applications of a single layer under its equivalent single exponential is illustrative.
The leading order term after the sum of the driving and mixing Hamiltonians is their single commutator
$-\frac{\gamma_{j}*\theta_{j}}{2}[H_{m},H_{c}]=-\frac{i\gamma_{j}*\theta_{j}}{2}\sum_{k,l\in E}\omega_{kl}(Y_{k}Z_{l}+Z_{k}Y_{l})\,.$ (S4)
Previous work in accelerating convergence within QAOA found these are precisely the terms most important for accelerating convergence Chandarana et al. (2022b); Zhu et al. (2022b).
Is it clear from the following relation
$\exp(\frac{i\pi ZY}{2})\left\{\begin{array}[]{lr}\left|00\right\rangle\pm\left|11\right\rangle\\ \left|01\right\rangle\pm\left|10\right\rangle\end{array}\right\}=\left\{\begin{array}[]{lr}\mp\left|01\right\rangle\mp\left|10\right\rangle\\ \quad\left|00\right\rangle\pm\left|11\right\rangle\end{array}\right\}$ (S9)
that $ZY$ rotations exactly rotate norm between the ±$\pm 1$ sector of $ZZ$ terms of the Hamiltonian. Based on these findings, we employ a parameterized ansatz of the form
$\left|\Psi(\theta)\right\rangle=\prod_{j=\{(r,q):r<q\}\in|V|}\exp(i\theta_{j}Z_{r}Y_{q})\left|+\right\rangle^{\otimes}$ (S10)
where $j$ indexes all unique pairs of vertices. The ordering is important for robust convergence, so we introduce the quantity
$\rho_{j}=\sum_{k\in E}|\omega_{jk}|$ (S11)
which represents the relative importance of each vertex in the overall $H_{c}$. Thus in this work, we order the $ZY$ rotations in the ansatz by their corresponding size in $\rho$. For a vertex set $|V|$ indexed now by their order in $\rho$, the ansatz exploited in this work has an explicit form
$\left|\Psi(\theta)\right\rangle=\exp(i\theta_{j}Z_{N-1}Y_{N})\ldots\exp(i\theta_{j}Z_{0}Y_{2})\exp(i\theta_{j}Z_{0}Y_{1})\left|+\right\rangle^{\otimes}\,.$ (S12)
Figure S1: Average error convergence for 100 random instances of (a) 3 regular, (b) NWS, and (c) SK graphs with 16 vertices and relaxed rounding performed. The orange line is VAR-IT, the green line is (3) optimized with ADAM, and the purple line shows the optimized error from 8-layer QAOA [from Dupont and Sundar (2024); Dupont (2023)].
IV Problem types
The different problem types explored in this work are dictated by the $\omega_{ij}$ between vertices, commonly referred to as the adjacency matrix. All graphs in this work are generated by the Python package NetworkX Hagberg et al. (2008).
Sherrington-Kirkpatrick (SK) – The SK model in this work is defined as the fully connected graph type, with $\omega_{ij}$ chosen between ±$\pm 1$ chosen with equal probability.
Newman-Watts-Strogatz small-world graphs (NWS) – In order to create models of this type, vertices of a ring graph are connected with its $k$ nearest neighbors. In this work, $k=4$. Afterwards, for each resulting edge (i,j), in the ring with connections to k nearest neighbors, add a new edge (i,l) with probability $p$, in this work $p=0.5$, to a randomly-chosen vertex $l$. Then $\omega_{ij}$ are assigned randomly in the interval $(0,1]$.
Random 3 Regular (3Reg) – This model arises from graphs where each vertex is connected to three other vertices randomly. $\omega_{ij}$ are all assigned a uniform value of 1 in this case.
V Relaxed Rounding
In Dupont and Sundar (2024), the authors introduced a quantum analogue of the classical relax and round method. They showed that it can easily be used to augment wavefunction methods like QAOA, and always improves the approximation ratio. It relies on the establishing the correlation matrix $\chi_{ij}$, defined as
$\chi_{ij}=(\delta_{ij}-1)\langle Z_{i}Z_{j}\rangle\,.$ (S13)
Then $\chi$ can be diagonalized classically to yield its eigenvectors $\{\alpha_{\nu}\in\mathbb{R}^{N}\}$. These can be translated to their closest bitstrings by rounding $\{z_{\nu}\leftarrow sign(\alpha_{\nu})\in{\pm 1}^{N}\}$. Then the $z_{\nu}$ which minimizes the $C(z)$ is chosen as the outcome of this quantum,relax, and round procedure. We apply this procedure at each iteration of our optimization procedure and report the findings in Fig. S1.
VI Further Numerical Results
In numerical results section of this work, we presented the error of the approximation ratio when using our VAR-IT method for optimization of Eq. 3, and the ADAM optimizer.
Figure S2: Average iterations required as a function of graph size.
Figure S3: Fraction of instances with $<1\%$ norm on optimal solutions after 100 iterations.
In addition to the VAR-IT method always converging to an optimal solution, Fig. S2 demonstrates that for the sizes explored in this work, the iterations required to reach convergence for all three problem types explored in this work appear to be scaling linearly in problem size. Additionally, Fig. S3 shows that an ansatz optimized with ADAM will yield increasing failures as problem size grows, while there does not appear to be any such signature with the VAR-IT optimization procedure.